OpenStudy (anonymous):

Solve 4x2 = 20x x = 4, x = 5 x = 4, x = −5 x = 0, x = 5 x = 0, x = −5 When the quadratic equation y = (x − 2)(3x + 1) is written in standard form, what is the value of the coefficient "b"? −5 3 −2 5 What are the x-intercept(s) of the graph of y − x2 = −3x − 10? (2, 0) and (−5, 0) (−2, 0) and (5, 0) (−2, 0) and (−5, 0) (2, 0) and (5, 0) Identify which equation is a quadratic equation. y = −5x − 3 y = −4x(x + 10) + 34 y + x2 = (x − 4)(x + 3) y − 2 Determine whether the graph of y = 2x2 − 5x + 8 opens up or down and whether it has a maximum or minimum point. Down and Maximum Down and Minimum Up and Minimum Up and Maximum What are the solutions of 6x2 − x − 15 = 0? x = 5 over 6, x = −3 x = −5 over 6, x = 3 x = 3 over 2, x = −5 over 3 x = −3 over 2, x =5 over 3 Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points) Please explain how to solve this. I want to learn! :)

OpenStudy (anonymous):

sorry you dont have to answer all just need help with a few :)

OpenStudy (anonymous):

I would love to understand how to do this :(

OpenStudy (anonymous):

4x2=20x 4x2-20x=0 4x(x-5)=0 either 4x=0 or x-5=0 or both. so x=0 and x=5.

OpenStudy (anonymous):

y=(x-2)(3x+1) y=\[3x ^{2}+x-6x-2\] y=\3x ^{2}-5x-2\ standard form of quad equation is ax2+bx+c=0 so coeff of b is -5

OpenStudy (anonymous):

y=-4x(x+10+34) is only the quad eqn of all as it has x raised to power 2.no one else has. in option 3 power of 2 get cancelled.

OpenStudy (anonymous):

awesome! I'm understanding it thank you so much! :)

OpenStudy (anonymous):

6x2-x-15=0 6x2-10x+9x-15=0 2x(3x-5)+3(3x-5)=0 (2x+3)(3x-5)=0 either of these are 0 or both so 2x+3=0 2x=-3 x=-3/2 and 3x-5=0 3x=5 x=5/3 so x=-3over 2 and 5over3

OpenStudy (anonymous):

awesome!! Thank you 100,00 times your a really big help :)

OpenStudy (anonymous):

anytime dear:) just in case m not here mail ur queries at rogerspoj@gmail.com

OpenStudy (anonymous):

Great I will definitely do that!

OpenStudy (anonymous):

2x^2-5x+8=0 has a min point at(1.25,4.875) and has shape like bowl. http://graph.tk/

OpenStudy (anonymous):

ok :)